OpenStudy (anonymous):

2.The volume of a rectangular prism is 183 in³. What is the volume of a pyramid which has the same height and a congruent base?

OpenStudy (lgbasallote):

\[\Large \mathsf{\text{Volume of rectangular prism = base}} \times \mathsf{\text{height}}\] \[\large \mathsf{\text{Volume of pyramid with rectangular base =}} \frac{1}{3} \mathsf{base} \times \mathsf{height}\]

OpenStudy (lgbasallote):

see a pattern?

OpenStudy (anonymous):

nope

OpenStudy (lgbasallote):

the volume of the pyramid (with rectangular base) is 1/3 of the volume of the rectangular prism